# How Does Quantization Affect Multilingual LLMs?

### v1

@@ Abstract
Quantization techniques are widely used to improve inference speed and deployment of large language models. While a wide body of work examines the impact of quantized LLMs on English tasks, none have examined the effect of quantization across languages. We conduct a thorough analysis of quantized multilingual LLMs, focusing on their performance across languages and at varying scales. We use automatic benchmarks, LLM-as-a-Judge methods, and human evaluation, finding that (1) harmful effects of quantization are apparent in human evaluation, and automatic metrics severely underestimate the detriment: a 1.7% average drop in Japanese across automatic tasks corresponds to a 16.0% drop reported by human evaluators on realistic prompts; (2) languages are disparately affected by quantization, with non-Latin script languages impacted worst; and (3) challenging tasks such as mathematical reasoning degrade fastest. As the ability to serve low-compute models is critical for wide global adoption of NLP technologies, our results urge consideration of multilingual performance as a key evaluation criterion for efficient models.
@@ 1 Introduction
With the shift towards large models, quantization is a widely adopted technique to reduce cost, improve inference speed, and enable wider deployment of LLMs. Work on quantization, however, is by-and-large evaluated in English only (e.g. Xiao et al., 2023; Ahmadian et al., 2024; Frantar et al., 2022). No works to our knowledge have characterized the impact of quantization on the multilingual generation capabilities expected from modern LLMs. Ubiquitous use of compression techniques in the real world drives urgency to the question how are multilingual models impacted?
@@
We analyze four state-of-the-art multilingual LLMs across 3 different sizes ranging from 8 to 103 billion parameters and covering up to 23 languages, under various quantization techniques. Critically, it is vital that we move beyond automatic evaluation and gather real human feedback on performance cost. We thus perform multilingual human evaluation on challenging real-world prompts in addition to LLM-as-a-Judge and evaluation on standard automatic benchmarks such as multilingual MMLU Hendrycks et al. (2020), MGSM Shi et al. (2023), and FLORES-200 (Costa-jussà et al., 2022a). Across experimental set-ups we find that:
@@
    Automatic metrics severely underestimate damage from quantization. Automatic evaluations estimate performance deterioration relative to FP16 across tasks at $-0.3\%$ (French) and $-1.7\%$ (Japanese) vs $-16.6\%$ and $-16.0\%$ reported by human evaluators.
@@
    Quantization affects languages differently. Non-Latin script languages are more greatly harmed on average. Across tasks, Latin-script languages scored $-0.7\%$ relative to FP16 for a 103B parameter model while non-Latin scripts scored $-1.9\%$. For a smaller 8-billion parameter model, scores were $-3.0\%$ vs. $-3.7\%$.
@@
    Occasionally, quantization brings benefits. Similar to Ahia et al. (2021) and Ogueji et al. (2022) on sparsity, we find that quantization benefits model performance in some cases: e.g., an average 1.3% boost across tasks for a 35B model quantized with W8A8.
@@ 2 Background
$\mathbf{W}_{Q}=\Delta\cdot\left\lfloor\frac{\mathbf{W}}{\Delta}\right\rceil,\quad\Delta=\frac{\max(|\mathbf{W}|)}{2^{N-1}}$ (1)
@@
${\mathbf{W}_{Q}}_{:,j}=\left\lfloor\frac{\mathbf{W}_{:,j}}{\Delta^{W}_{:,j}}\right\rceil,\Delta^{W}_{:,j}=\frac{\max(|\mathbf{W}_{:,j}|)}{2^{N-1}}$ (2)
${\mathbf{X}_{Q}}_{i,:}=\left\lfloor\frac{\mathbf{X}_{i,:}}{\Delta^{X}_{i,:}}\right\rceil,\Delta^{X}_{i,:}=\frac{\max(|\mathbf{X}_{i,:}|)}{2^{N-1}}$ (3)
$Y=\Delta^{X}\odot(\mathbf{X}_{Q}\mathbf{W}_{Q})\odot\Delta^{W}$ (4)
where, $\Delta^{W}\in\mathbb{R}^{d_{out}}$, $\Delta^{X}\in\mathbb{R}^{seq}$ and $\odot$ denotes element-wise multiplication by broadcasting the elements to match the shape of the operands.
3 Experiment Set-up
@@
We evaluate Command R+ [^1] https://docs.cohere.com/docs/command-r-plus, Command R [^2] https://docs.cohere.com/docs/command-r, and Aya 23 models (Aryabumi et al., 2024) as representatives of state-of-the-art multilingual LLMs. Command models are 103 and 35 billion parameters, and Aya 23 models are 35 and 8 billion parameters. We quantize the models using the weights available on HuggingFace.
@@
For Command R and R+, we evaluate both weight-only quantization at 8-bit (W8 with per-column scaling) and 4-bit (W4-g with group-wise scaling using GPTQ (Frantar et al., 2022)), as well as weight-and-activation quantization at 8-bit (W8A8 with per-column scaling for weights and per-token scaling for activations).
Ahmadian et al. (2024) show that if the model is trained with the right hyper-parameters, naive Weight-and-Activation Quantization has minimal degradation. Otherwise, one may leverage SmoothQuant (Xiao et al., 2023) to smoothen the distribution of activations making them more amenable to quantization. We therefore also explore W8A8-SmoothQuant (a W8A8 variant with SmoothQuant) for Command R+ as well as a 4-bit weight-only quantized variant with column-wise scaling (W4) to understand the impact of scaling granularity at extremely low-bit precision. Following (Frantar et al., 2022; Xiao et al., 2023), we use 128 English samples for calibration for SmoothQuant and GPTQ.
For Aya 23 8B and 35B, we use bitsandbytes [^3] https://github.com/TimDettmers/bitsandbytes to obtain 8-bit and 4-bit quantized models. Bitsandbytes uses LLM.int8() (Dettmers et al., 2022)—similar to W8A8 described above except that it performs certain computations in FP16. Bitsandbytes 4-bit uses the NF4 datatype (Dettmers et al., 2023) to perform Quantile Quantization which limits degradation at the expense of inference speedups.
@@ 3.1 Automatic Evaluation
Raw numeric results are in the Appendix. Results are averaged over 5 runs.444k=0, p=0.75, temp=0.3, except mMMLU, which, as a QA eval, is run deterministically with temp=0.
Multilingual MMLU (mMMLU)
This multi-domain question answering task consists of 14,000+ multiple-choice questions. We translate MMLU (Hendrycks et al., 2020) to 9 languages with Google Translate and refer to this version as mMMLU. We measure accuracy in a 5-shot setting. An example is in Table A1.
@@
MGSM is a generative mathematics evaluation set manually translated from GSM8K (Cobbe et al., 2021). Of our target languages, it is available for German, Spanish, French, Japanese, and Chinese. We report accuracy over the 250-item test set for each language.
@@
Aya 23 models are evaluated using an extended version of the Aya evaluation setup (Aryabumi et al., 2024) using the unseen discriminative tasks (XWinograd (Muennighoff et al., 2023), XCOPA (Ponti et al., 2020), XStoryCloze (Lin et al., 2022)), mMMLU (Okapi; Dac Lai et al., 2023), MGSM, and Belebele (Bandarkar et al., 2023) from eval-harness (Gao et al., 2023). [^6] We follow the setup used by Üstün et al. (2024): each evaluation is run once, and for FLORES, no sampling is used and metric is spBLEU. We evaluate models on languages included in the covered 23 languages, except for the unseen tasks where we use all available languages.777mMMLU: ar, de, es, fr, hi, id, it, nl, pt, ro, ru, uk, vi, zh. MGSM: de, es, fr, ja, ru, zh. Belebele: {mMMLU} + cs, fa, el, ja, ko, pl, tr. FLORES: {Belebele} + he. Aya evaluations allow us to add: Czech, Greek, Hebrew, Hindi, Indonesian, Dutch, Persian, Polish, Romanian, Russian, Turkish, Ukrainian, Vietnamese.
@@ 3.2 Human Evaluation
We run human evaluation in Spanish, French, Korean, and Japanese.
@@
Annotator Statistics
Annotations and translations were completed by native-level speakers of the respective languages, each of whom is also fluent in English. Annotators were paid by the hour, with compensation above the federal minimum wage of the country of employment.
Annotation Interface
We use a pairwise evaluation setup. Annotators see a prompt and two (shuffled) completions of the FP16 model and a quantized variant. They rate each response on a 5-point Likert scale, then express a preference between the two model outputs (tie, weak preference, strong preference). We encourage annotators to avoid tied rankings. Win rates are based on the ranking preferences alone.
3.3 LLM/RM-as-a-Judge
Because human evaluation is costly and time-intensive, it is common to use an “LLM-as-a-Judge” to rate model completions (e.g. Li et al., 2023b; Zheng et al., 2023). Reward models (RMs) can also simulate human preference. An RM scores multiple completions given the same prompt, and the prompt-completion pair with the higher score is deemed preferred. We call this RM-as-a-Judge.
We assess quantized model outputs using LLM- and RM-as-a-Judge. In the former, an LLM selects a preferred response from a $<$instruction, modelA_completion, modelB_completion$>$ tuple (see Table A2). Following Üstün et al. (2024); Aryabumi et al. (2024) we use GPT-4888gpt-4-turbo (gpt-4-1106-preview): https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 as an LLM proxy judge. To minimize bias, we randomize the order of model outputs. For RM-as-a-Judge, a multilingual reward model scores $<$prompt, completion$>$ pairs for each model output, over which we calculate win-rate. We report win-rates of quantized models versus the FP16 baseline.
4 Results
@@ 4.1 By Quantization Level
Command R and R+
In Table 1, we aggregate results of each metric for each level of quantization. We average scores across languages, then calculate the relative percentage drop from FP16.999Ex. For 103B W4-g MGSM, scores were: {de: 71.2, es: 75.7, fr: 69.0, ja: 58.0, zh: 68.9}, thus the average score was 68.6—a 2.9% drop from FP16 ($\frac{68.6-70.6}{70.6}=-0.029$). We discuss results of W8, W8A8, and W4-g quantization, which are variants available for both Command model sizes. Most results follow intuition: greater quantization leads to larger performance degradation: $-0.2\%$ for W8, $-0.8\%$ for W8A8, and $-0.9\%$ overall for the 103B model. An exception is W8A8 for the 35B, which experiences a slight boost overall due to higher performance on translation and language confusion evaluations.
Aya 23 Models
4.2 By Task
Are tasks differently affected by quantization?
Results here reference Tables 1 and 2, with full raw and relative results in Appendix A.2. Mathematical reasoning as measured by MGSM is strikingly affected by quantization. Relative performance of the 35B W4-g model is a dismal $-13.1\%$ on average, with as poor as $-17.3\%$ in Chinese. MGSM and Belebele are most greatly degraded for Aya 23 models with W4 quantization, dropping $7.5\%$ and $8.5\%$ on the 8B model. mMMLU is the next most greatly degraded task.
@@ 4.2 By Task
Curiously, there are some fleeting performance boosts: an increase of 1.8–2.1% on MGSM and mild improvements on FLORES with W8 on Aya models, and a similar translation boost of the 35B Command model at W8A8. Quantization generally has no effect or causes mild improvement on the monolingual language confusion task, and cross-lingual language confusion performance is boosted with greater quantization in some cases.
@@ 4.3 By Language
W4-g causes considerable degradation across languages for the 35B Command model. A relationship between language and performance is apparent: $\neg$¬Ltn/IE languages typically degrade more. Chinese, Japanese, and Korean are particularly harmed by W4 quantization of the 103B. The effect is seen consistently across all automatic metrics for Command, with limited exception. Table 5 is discussed more thoroughly in Section 4.5, but also shows this discrepancy. In the Appendix, we see the same for Aya 23 models at W4.

### v2

@@ Abstract
Quantization techniques are widely used to improve inference speed and deployment of large language models. While a wide body of work examines the impact of quantization on LLMs in English, none have evaluated across languages. We conduct a thorough analysis of quantized multilingual LLMs, focusing on performance across languages and at varying scales. We use automatic benchmarks, LLM-as-a-Judge, and human evaluation, finding that (1) harmful effects of quantization are apparent in human evaluation, which automatic metrics severely underestimate: a 1.7% average drop in Japanese across automatic tasks corresponds to a 16.0% drop reported by human evaluators on realistic prompts; (2) languages are disparately affected by quantization, with non-Latin script languages impacted worst; and (3) challenging tasks like mathematical reasoning degrade fastest. As the ability to serve low-compute models is critical for wide global adoption of NLP technologies, our results urge consideration of multilingual performance as a key evaluation criterion for efficient models.
@@ 1 Introduction
With the shift towards large models, quantization is a widely adopted technique to reduce cost, improve inference speed, and enable wider deployment of LLMs. Work on quantization, however, is by-and-large evaluated in English only (e.g. Xiao et al., 2023; Ahmadian et al., 2024; Frantar et al., 2022).
Figure 1: Automatic metrics severely underestimate damage from quantization. Shown: 103B W4 quantized Command model with group-wise scaling vs. FP16. Avg: mMMLU, FLORES, Language Confusion (LC). English avg: mMMLU, MGSM, monolingual LC.
No works to our knowledge have characterized the impact of quantization on the multilingual generation capabilities expected from modern LLMs. Ubiquitous use of compression techniques in the real world drives urgency to the question how are multilingual models impacted?
@@
We analyze four state-of-the-art (SOTA) multilingual LLMs across 3 different sizes ranging from 8 to 103 billion parameters and covering up to 23 languages, under various quantization techniques. Critically, it is vital that we move beyond automatic evaluation and gather real human feedback on performance cost. We thus perform multilingual human evaluation on challenging real-world prompts in addition to LLM-as-a-Judge and evaluation on standard automatic benchmarks such as multilingual MMLU Hendrycks et al. (2020), MGSM Shi et al. (2023), and FLORES-200 (Costa-jussà et al., 2022a). Across experimental set-ups we find that:
@@
    Automatic metrics severely underestimate damage from quantization. Automatic evaluations estimate deterioration relative to FP16 across tasks at $-0.3\%$ (French) and $-1.7\%$ (Japanese) vs. $-16.6\%$ and $-16.0\%$ reported by human evaluators. See Figure 1.111Figure excludes MGSM (not available for Korean.)
@@
    Quantization affects languages differently. Degradation on automatic metrics appears negatively correlated with training data set size, and non-Latin script languages are more harmed on average. Across tasks, Latin-script languages scored $-0.7\%$ relative to FP16 for a 103B parameter model while non-Latin scripts scored $-1.9\%$. For a smaller 8-billion parameter model, scores were $-3.0\%$ vs. $-3.7\%$.
@@
    Occasionally, quantization brings benefits. Similar to Badshah and Sajjad (2024)’s finding on English tasks, we find that quantization benefits multilingual model performance in some cases: e.g., an average 1.3% boost across tasks for a 35B model quantized with W8A8. This aligns with findings on the benefit of other compression methods such as sparsity (Ahia et al., 2021; Ogueji et al., 2022).
@@ 2 Background
$\mathbf{W}_{Q}=\left\lfloor\frac{\mathbf{W}}{\Delta}\right\rceil,\quad\Delta=\frac{\max(|\mathbf{W}|)}{2^{N-1}-1}$ (1)
@@
${\mathbf{W}_{Q}}_{:,j}=\left\lfloor\frac{\mathbf{W}_{:,j}}{\Delta^{W}_{:,j}}\right\rceil,\Delta^{W}_{:,j}=\frac{\max(|\mathbf{W}_{:,j}|)}{2^{N-1}-1}$ (2)
${\mathbf{X}_{Q}}_{i,:}=\left\lfloor\frac{\mathbf{X}_{i,:}}{\Delta^{X}_{i,:}}\right\rceil,\Delta^{X}_{i,:}=\frac{\max(|\mathbf{X}_{i,:}|)}{2^{N-1}-1}$ (3)
where $\Delta^{W}\in\mathbb{R}^{d_{out}}$ and $\Delta^{X}\in\mathbb{R}^{seq}$. In the forward pass, $\mathbf{Y}$ is calculated as below, where $\odot$ denotes element-wise multiplication by broadcasting the elements to match the shape of the operands. The multiplication in lower-precision $\mathbf{X}_{Q}\mathbf{W}_{Q}$ is what leads to throughput gains. Multiplying by $\Delta^{W}$ and $\Delta^{X}$ de-quantizes the result so that $\mathbf{Y}$ has the same (higher) precision as the original $\mathbf{X}$.
$\mathbf{Y}=\Delta^{X}\odot(\mathbf{X}_{Q}\mathbf{W}_{Q})\odot\Delta^{W}$ (4)
3 Experimental Set-up
@@
We use Command R+ [^2] https://docs.cohere.com/docs/command-r-plus, Command R [^3] https://docs.cohere.com/docs/command-r, and Aya 23 models (Aryabumi et al., 2024) as representative of SOTA multilingual LLMs. Command models are 103 and 35 billion parameters (R+/R). Aya 23 models are 35 and 8 billion parameters. We quantize the weights available on HuggingFace.
@@
For Command R/R+ (35B/103B), we evaluate weight-only quantization at 8-bit (W8 with per-column scaling) and 4-bit (W4-g with group-wise scaling using GPTQ (Frantar et al., 2022)), as well as weight-and-activation quantization at 8-bit (W8A8 with per-column scaling for weights and per-token scaling for activations).
When trained with the right hyper-parameters, naive Weight-and-Activation Quantization has minimal degradation (Ahmadian et al., 2024). Otherwise, SmoothQuant (Xiao et al., 2023) may smoothen the activation distributions to be more amenable to quantization. We explore W8A8-SmoothQuant (W8A8 with SmoothQuant) for Command R+ (103B) and a 4-bit weight-only quantized variant with column-wise scaling (W4) to understand the impact of scaling granularity at extremely low-bit precision. We use 128 English samples for calibration for SmoothQuant and GPTQ (Frantar et al., 2022; Xiao et al., 2023).
For Aya 23 8B and 35B, we use bitsandbytes [^4] https://github.com/TimDettmers/bitsandbytes for 8-bit and 4-bit quantization. This uses LLM.int8() (Dettmers et al., 2022)—similar to W8A8 except it performs some computations in FP16. The 4-bit uses the NF4 datatype (Dettmers et al., 2023) to perform Quantile Quantization which limits degradation at the expense of inference speedups.
@@ 3.1 Automatic Evaluation
Raw numeric results are in the Appendix. Results are averaged over 5 runs.555k=0, p=0.75, temp=0.3, except mMMLU, which, as a QA eval, is run deterministically with temp=0.
Multilingual MMLU
14,000+ multi-domain multiple-choice questions. We translate MMLU (Hendrycks et al., 2020) to 9 languages with Google Translate and call it mMMLU. We measure 5-shot accuracy. Example in Table A1.
@@
Generative mathematics test set manually translated from GSM8K (Cobbe et al., 2021). Of our target languages, it is available for German, Spanish, French, Japanese, Chinese. We report accuracy for each language.
@@
Aya 23 models are evaluated using an extended version of the Aya evaluation setup (Aryabumi et al., 2024) using the unseen discriminative tasks—those where there is no dataset in the models’ training mixture from the same task categories (XWinograd (Muennighoff et al., 2023), XCOPA (Ponti et al., 2020), XStoryCloze (Lin et al., 2022)), mMMLU (Okapi; Dac Lai et al., 2023), MGSM, and Belebele (Bandarkar et al., 2023) from eval-harness (Gao et al., 2023). [^7] We follow Üstün et al. (2024): each evaluation is run once; For FLORES, no sampling is used and metric is spBLEU. We evaluate models on languages included in the covered 23 languages, except for the unseen tasks where we use all available languages.888mMMLU: ar, de, es, fr, hi, id, it, nl, pt, ro, ru, uk, vi, zh. MGSM: de, es, fr, ja, ru, zh. Belebele: {mMMLU} + cs, fa, el, ja, ko, pl, tr. FLORES: {Belebele} + he. Aya evaluations allow us to add: Czech, Greek, Hebrew, Hindi, Indonesian, Dutch, Persian, Polish, Romanian, Russian, Turkish, Ukrainian, Vietnamese.
@@ 3.2 Human Evaluation
We run human evaluation in Spanish, French, Korean, Japanese, and English.
@@
Annotation
Annotations and translations were completed by native-level speakers of the respective languages who are also fluent in English. [^9] Paid hourly, above min. wage of country of employment. The annotation interface supports pairwise evaluation. Annotators see a prompt and two (shuffled) completions of the FP16 model and a quantized variant which they rate on a 5-point Likert scale, then express a preference (tie, weak preference, strong preference). We encourage annotators to avoid ties. Win rates are based on ranking preferences alone.
@@
3.3 LLM/RM-as-a-Judge
Because human evaluation is costly and time-intensive, it is common to use an “LLM-as-a-Judge” to rate model completions (e.g. Li et al., 2023b; Zheng et al., 2023). Reward models (RMs) can also simulate human preference. A RM scores multiple completions given the same prompt, and the prompt-completion pair with the higher score is deemed preferred. We call this RM-as-a-Judge.
We assess quantized model outputs using LLM- and RM-as-a-Judge. In the former, an LLM selects a preferred response from a $<$instruction, modelA_completion, modelB_completion$>$ tuple (see Table A2). We use GPT-4101010turbo (gpt-4-1106-preview): https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 as an LLM proxy judge following Üstün et al. (2024) and Aryabumi et al. (2024). We randomize the order of model outputs to minimize bias. For RM-as-a-Judge, a multilingual RM scores $<$prompt, completion$>$ pairs for each model output, over which we calculate win-rate. We report win-rates of quantized models versus the FP16 baseline.
4 Results
@@ 4.1 By Quantization Level
Command R (35B) and R+ (103B)
In Table 1, we aggregate results of each metric for each level of quantization. We average scores across languages, then calculate the relative percentage drop from FP16.111111Ex. For 103B W4-g MGSM, scores were: {de: 71.2, es: 75.7, fr: 69.0, ja: 58.0, zh: 68.9}, thus the average score was 68.6—a 2.9% drop from FP16 ($\frac{68.6-70.6}{70.6}=-0.029$). We discuss results of W8, W8A8, and W4-g quantization, which are variants available for both Command model sizes. Most results follow intuition: greater quantization leads to larger performance degradation: $-0.2\%$ for W8, $-0.8\%$ for W8A8, and $-0.9\%$ for W4-g of the 103B model. An exception is W8A8 for the 35B which shows a slight boost overall due to higher performance on translation and language confusion evaluations.
@@
Aya 23 Models
4.2 By Task
Are tasks differently affected by quantization?
@@ 4.2 By Task
There are some fleeting performance boosts: +1.8–2.1% on MGSM and mild improvements on FLORES with W8 on Aya models, and a similar translation boost of the 35B Command model at W8A8. Quantization generally has no effect or causes mild improvement on the monolingual language confusion task, and cross-lingual language confusion performance is boosted with greater quantization in some cases.
@@ 4.3 By Language
W4-g causes considerable degradation across languages for the 35B Command model. A relationship with language is apparent: $\neg$¬Ltn/IE languages typically degrade more. Chinese, Japanese, and Korean are particularly harmed by W4 on the 103B. The effect is seen consistently across all automatic metrics for Command, with limited exception. Table 6 is discussed more thoroughly in Section 4.5, but also shows this discrepancy. In the Appendix, we see the same for Aya 23 models at W4.
@@
Avg Rel. %$\%\Delta$ mMMLU MGSM Lang. Conf. (Mono)
en Ltn/IE All en Ltn/IE All en Ltn/IE All en Ltn/IE All
W8 0.1% -0.3% -0.3% 0.0% 0.0% 0.0% 0.3% -0.7% -1.0% 0.0% -0.1% 0.0%
W8A8-sq -1.2% -0.6% -0.7% -0.1% -0.4% -0.5% -3.4% -1.3% -1.6% 0.0% -0.1% 0.0%
W8A8 -0.3% -0.7% -0.8% -0.7% -1.3% -1.7% 0.0% -0.9% -1.0% -0.1% 0.0% 0.2%
W4-g -2.0% -0.9% -1.5% -1.7% -1.1% -1.5% -4.4% -1.8% -3.0% 0.0% 0.1% 0.0%
103B W4 -3.7% -3.9% -4.3% -3.3% -3.9% -4.4% -7.9% -8.0% -8.8% 0.0% 0.1% 0.1%
W8 -0.1% -0.2% -0.2% -0.1% -0.1% -0.1% -0.3% -0.6% -0.8% 0.1% 0.1% 0.1%
W8A8 0.2% -1.6% -1.8% 0.0% 0.0% -0.2% 0.0% -4.9% -5.6% 0.7% 0.0% 0.3%
35B W4-g -1.2% -4.8% -5.2% -1.8% -1.5% -2.0% -2.2% -12.2% -13.1% 0.4% -0.6% -0.4%
Table 5: Relative performance of quantized Command models in English vs. other languages. All non-English languages (All), non-English Latin-script/Indo-European languages (Ltn/IE).
How does training data size affect performance?
Training mixtures for Command and Aya 23 models are not released, so a definitive relationship between data set size and downstream performance cannot be determined. Instead, we might assume the training data follows the distribution from well-known large multilingual corpora. In Figure 2, we correlate per-language relative performance vs. FP16 from Table 3 with amount of data in mC4 (Xue et al., 2021). [^13] https://github.com/allenai/allennlp/discussions/5265. Correlation with size in tokens from Xue et al. (2021)’s Table 6 shows similar $R^{2}$. Data size by lang. (GB): [ar: 57, de: 347, es: 433, fr: 318, it: 162, ja: 164, ko: 26, pt: 146, zh: 39]. English excluded as it cannot be averaged with FLORES & cross-lingual Language Confusion. The correlation between downstream performance and data set size is stronger as quantization becomes more extreme: from $R^{2}=0.24$ for W8 to $R^{2}=0.63$ with W4.
Figure 2: Data size in mC4 (Xue et al., 2021) vs. avg. perf. under quantization. Table 3, Command 103B.
How does performance compare to English?
Table 5 shows relative performance of quantized Command 103B and 35B models in English vs. other languages for tasks which could be evaluated in English. [^14] FLORES / cross-lingual Language Confusion cannot be. Under most settings, English degrades less than the average of all others. The largest gap is on MGSM for the 35B, where the model is very sensitive to W8A8 and W4-g quantization outside of English. Results for the Aya 23 models are in Table A17, where performance is worse on average for non-English languages at W4, while being less consistent at W8.